# Breaking the Dimensional Barrier: Dynamic Portfolio Choice with Parameter Uncertainty via Pontryagin Projection

### v2

@@ 4.2 High-dimensional CRRA benchmark: projection and amortization
Mechanism: why misalignment can reduce projection gains. To explain when and why the projection gains shrink, we analyze Stage 2 diagnostic statistics reported in Appendix G; see Figures 47. The diagnostics indicate that the degradation under misalignment is driven primarily by increased stationarity residuals and curvature mismatch, rather than by catastrophic denominator sign failures: (i) the Stage 2 residual norm grows with dimension and becomes especially large in the hardest misaligned regime, (ii) the projection denominator magnitude stays away from zero at typical quantiles, and (iii) the bad-sign fraction remains negligible, while (iv) the effective curvature statistic $\kappa$ stays near the nominal $1/\gamma$ reference in easy regimes but can deviate substantially in the hardest misaligned/high-uncertainty setting. These patterns are consistent with the geometric explanation: when $P$ and $\Sigma$ do not commute, the inverse operations implicit in projection mix uncertainty directions that are poorly aligned with risk directions, which can amplify estimation noise in the projected correction as $d$ grows.
@@
Pathwise sanity check. Figure 3 complements the decision-time RMSE with a trajectory-level view under common random numbers. In the aligned case, the on-the-fly Stage 2 teacher tracks the analytic $q$-reference closely along a realized path and reduces the deviation $\Delta\log X_{t}$ relative to the warm Stage 1 policy. In the misaligned case, the teacher can deviate more noticeably under the same common-noise protocol, mirroring the reduced projection advantage in the hardest regimes of Figure 2 and motivating amortization/reliability mechanisms in interactive distillation.
@@ 4.3 A strong RL baseline: PPO, and why it falls short in our benchmark
Since classical HJB solvers and value-function-based deep PDE surrogates are not practical baselines in the high-dimensional uncertain regime targeted here (Section B), PPO serves as a strong simulation-only comparator that operates on the same sampled trajectories without exploiting value-function PDE structure. For a fair comparison, PPO is trained on the same Euler simulator and time discretization as our PG–DPO pipeline, under the same deployability restriction (the policy never observes the latent $\theta$), and under the same terminal-utility objective. We match simulation budgets at the level of total environment interactions (simulated trajectories $\times$ time steps); implementation and PPO tuning details are deferred to the appendix and code release.
Empirical outcome. Table LABEL:tab:crra_stage12_distill_ppo_landscape shows that PPO remains far from the analytic decision-time $q$-reference across essentially all conditions, with RMSE typically on the order of $10^{-1}$. In contrast, the Pontryagin-based pipeline attains substantially smaller errors: in aligned regimes Stage 2 projection reaches the $10^{-5}$$10^{-4}$ range for small and moderate uncertainty, while in misaligned regimes the projection advantage narrows but remains systematic. Distillation improves the deployable Stage 1 policy relative to basic PG–DPO, but does not eliminate the remaining gap to the post-hoc projection, consistent with the amortization interpretation in Section 4.2.
@@
Terminal-reward credit assignment. With terminal utility as the only reward signal, PPO faces a long-horizon credit-assignment problem: most intermediate transitions carry zero reward, so advantage estimates inherit high variance and weak learning signals, especially in higher dimensions.
@@
In contrast, Stage 1 exploits backpropagation through the differentiable simulator (pathwise gradients), and Stage 2 leverages the affine-in-control Pontryagin structure through a $q$-aggregated projection, replacing a noisy high-dimensional policy-gradient update by a structured stationarity correction that is tailored to the $\theta$-blind ex–ante objective.
@@ 5 Recovering Intertemporal Hedging Demand in Factor-Driven Markets
Sections 4 stressed scaling under static drift uncertainty, where the target $q$-reference is time-homogeneous and largely myopic. Here we shift the focus to an economic target: recovering the intertemporal hedging demand induced by factor-driven investment opportunities when return shocks are correlated with factor shocks (Campbell and Viceira, 2002; Xia, 2001).
We adopt the mean-reverting Gaussian premium benchmark of Section 2.3.2. At decision time ($t=0$), the investor is assumed to have access only to an interval/posterior uncertainty description $q(dy)$ for the initial premium state $Y_{0}$ (e.g., $Y_{0}\sim\mathcal{N}(m_{0},P_{0})$), and we restrict attention to deployable $Y$-blind policies that depend only on observable wealth and time-to-go. When return and factor shocks are correlated ($\rho\neq 0$), the analytic OU decision-time reference (43) contains the cross term $M_{\mathrm{cross}}(\tau)$ (with $\tau:=T-t$) that encodes intertemporal hedging demand.
We evaluate the two-stage PG–DPO / $q$-aggregated Pontryagin projection pipeline (and interactive distillation) against a model-free PPO baseline, all under the same $Y$-blind deployability restriction. Performance is measured by decision-time RMSE at $t=0$ relative to the analytic constant-portfolio OU reference (43).
@@ 5.1 Experimental setting
Decision-time interval/posterior estimate for $Y_{0}$ and deployability. Throughout this section, all reported policies are deployable and $Y$-blind: the control is a function of observable wealth and time-to-go only, and never takes as input (i) a point estimate of $Y_{0}$, (ii) the realized initial premium $Y_{0}$, or (iii) the factor path $(Y_{t})$ (including the PPO baseline). At deployment, the investor is assumed to have access only to an uncertainty description $q(dy)$ for $Y_{0}$ (interval/posterior output of an external estimation pipeline), and commits to a single policy optimized ex ante under $Y_{0}\sim q$ (without filtering/updating over $[0,T]$).
@@
Two-stage solver, amortization, and evaluation protocol. We use the two-stage pipeline of Section 3. Stage 1 trains a deployable policy by stochastic gradient ascent using pathwise/BPTT gradients (Section 3.1). Stage 2 applies the $q$-aggregated Pontryagin projection computed under a warm-up policy (Section 3.2), implemented in the residual/control-variate form (Section F.1). Interactive distillation amortizes the projected teacher into a fast deployable policy network (Section F.2). As a model-free baseline, we also train a PPO policy under the same $Y$-blind observation restriction and report its decision-time full RMSE in Table LABEL:tab:ou_rmse_s0_sweep_landscape.
@@
In addition to the full allocation error, we use the myopic+hedging decomposition induced by the OU structure, and report component-wise diagnostics for the projected (Stage 2) rules. To reduce noise from stochastic optimization, for each condition we summarize each metric by a tail median over the last six evaluation checkpoints.
@@ 5.2 Results: hedging-demand recovery, amortization, and robustness to decision-time uncertainty
Metrics and diagnostics. Table LABEL:tab:ou_rmse_s0_sweep_landscape reports the full decision-time RMSE at $t=0$ for all deployable policies, including the PPO baseline. To isolate the economic hedging channel, we use the OU-induced decomposition of the analytic reference
@@
where $\pi^{\mathrm{myo}}$ is the independence-case reference obtained by setting $\rho=0$ (equivalently, dropping $M_{\mathrm{cross}}(\tau)$ in (43)), and $\pi^{\mathrm{hedge}}:=\pi^{\mathrm{ref}}-\pi^{\mathrm{myo}}$ is the residual $\rho$-dependent term. We report the hedging-component RMSE for the projected (Stage 2) rules in Table 1; the myopic-component RMSE and hedging-direction cosine similarity are deferred to Appendix H. Since Stage 1 and PPO do not expose a compatible myopic/hedging split under our diagnostic protocol, component-wise diagnostics are reported only for Stage 2.
$s_{0}$ Method $d=5$ $10$ $50$ $100$
Aligned $P_{0}$
$10^{-3}$ Stage 1+Stage 2 (Basic) $4.59\text{\times}{10}^{-05}$ $4.87\text{\times}{10}^{-05}$ $1.37\text{\times}{10}^{-04}$ $1.55\text{\times}{10}^{-04}$
Stage 1+Stage 2 (Distill.) $4.39\text{\times}{10}^{-05}$ $5.25\text{\times}{10}^{-05}$ $1.37\text{\times}{10}^{-04}$ $1.42\text{\times}{10}^{-04}$
$10^{-2}$ Stage 1+Stage 2 (Basic) $98\text{\times}{10}^{-05} 4.4$$47\text{\times}{10}^{-05}$ $1.44\text{\times}{10}^{-04}$ $1.74\text{\times}{10}^{-04}$
Stage 1+Stage 2 (Distill.) $4.43\text{\times}{10}^{-05}$ $4.95\text{\times}{10}^{-05}$ $1.47\text{\times}{10}^{-04}$ $1.55\text{\times}{10}^{-04}$
$10^{-1}$ Stage 1+Stage 2 (Basic) $5.27\text{\times}{10}^{-05}$ $3.99\text{\times}{10}^{-05}$ $2.60\text{\times}{10}^{-04}$ $2.95\text{\times}{10}^{-04}$
Stage 1+Stage 2 (Distill.) $4.72\text{\times}{10}^{-05}$ $5.36\text{\times}{10}^{-05}$ $2.58\text{\times}{10}^{-04}$ $2.86\text{\times}{10}^{-04}$
Misaligned $P_{0}$
$10^{-3}$ Stage 1+Stage 2 (Basic) $5.00\text{\times}{10}^{-05}$ $4.87\text{\times}{10}^{-05}$ $1.34\text{\times}{10}^{-04}$ $1.57\text{\times}{10}^{-04}$
Stage 1+Stage 2 (Distill.) $4.30\text{\times}{10}^{-05}$ $5.33\text{\times}{10}^{-05}$ $1.40\text{\times}{10}^{-04}$ $1.43\text{\times}{10}^{-04}$
$10^{-2}$ Stage 1+Stage 2 (Basic) $4.93\text{\times}{10}^{-05}$ $5.65\text{\times}{10}^{-05}$ $1.53\text{\times}{10}^{-04}$ $1.75\text{\times}{10}^{-04}$
Stage 1+Stage 2 (Distill.) $4.56\text{\times}{10}^{-05}$ $6.09\text{\times}{10}^{-05}$ $1.53\text{\times}{10}^{-04}$ $1.56\text{\times}{10}^{-04}$
$10^{-1}$ Stage 1+Stage 2 (Basic) $5.45\text{\times}{10}^{-05}$ $1.55\text{\times}{10}^{-04}$ $3.19\text{\times}{10}^{-04}$ $3.28\text{\times}{10}^{-04}$
Stage 1+Stage 2 (Distill.) $5.20\text{\times}{10}^{-05}$ $1.57\text{\times}{10}^{-04}$ $3.13\text{\times}{10}^{-04}$ $3.20\text{\times}{10}^{-04}$
Table 1: Decision-time RMSE at $t=0$ for the hedging component of the OU decision-time reference, evaluated on the post-hoc projected (Stage 2) policies (tail median over the last six evaluation checkpoints). We use the decomposition $\pi^{\mathrm{ref}}(\tau)=\pi^{\mathrm{myo}}(\tau)+\pi^{\mathrm{hedge}}(\tau)$ where $\pi^{\mathrm{myo}}$ is the $\rho=0$ reference and $\pi^{\mathrm{hedge}}$ is the residual $\rho$-dependent term (cf. (43)). Component-wise diagnostics are reported for Stage 2 since Stage 1 and PPO do not explicitly output a compatible myopic/hedging split under our protocol.

### v3

@@ 4.2 High-dimensional CRRA benchmark: projection and amortization
Mechanism: why misalignment can reduce projection gains. To explain when and why the projection gains shrink, we analyze Stage 2 diagnostic statistics reported in Appendix H; see Figures 58. The diagnostics indicate that the degradation under misalignment is driven primarily by increased stationarity residuals and curvature mismatch, rather than by catastrophic denominator sign failures: (i) the Stage 2 residual norm grows with dimension and becomes especially large in the hardest misaligned regime, (ii) the projection denominator magnitude stays away from zero at typical quantiles, and (iii) the bad-sign fraction remains negligible, while (iv) the effective curvature statistic $\kappa_{\mathrm{curv}}$ stays near the nominal $1/\gamma$ reference in easy regimes but can deviate substantially in the hardest misaligned/high-uncertainty setting. These patterns are consistent with a geometric explanation: when $P$ and $\Sigma$ do not commute, the inverse operations implicit in the projection mix directions of uncertainty that are misaligned with risk directions, thereby amplifying estimation noise in the projected correction as $d$ grows.
@@
Pathwise sanity check. Figure 3 complements the decision-time RMSE with a trajectory-level view under common random numbers. In the aligned case, the on-the-fly Stage 2 teacher tracks the analytic $q$-reference closely along a realized path and reduces the deviation $\Delta\log X_{t}$ relative to the warm Stage 1 (DPO) policy. In the misaligned case, the teacher can deviate more noticeably under the same common-noise protocol, mirroring the reduced projection advantage in the hardest regimes of Figure 2 and motivating amortization/reliability mechanisms in interactive distillation.
Aligned
$s$ Method $d=5$ $10$ $50$ $100$ $10^{-3}$ DPO (basic) $8.76\text{\times}{10}^{-03}$ $3.74\text{\times}{10}^{-03}$ $1.81\text{\times}{10}^{-03}$ $2.58\text{\times}{10}^{-03}$ $10^{-3}$ PG–DPO (basic) $4.95\text{\times}{10}^{-05}$ $1.10\text{\times}{10}^{-05}$ $8.50\text{\times}{10}^{-06}$ $5.00\text{\times}{10}^{-06}$ $10^{-3}$ DPO (distill.) $4.10\text{\times}{10}^{-03}$ $2.48\text{\times}{10}^{-03}$ $1.31\text{\times}{10}^{-03}$ $1.70\text{\times}{10}^{-03}$ $10^{-3}$ PG–DPO (distill.) $5.55\text{\times}{10}^{-05}$ $1.10\text{\times}{10}^{-05}$ $8.00\text{\times}{10}^{-06}$ $5.50\text{\times}{10}^{-06}$ $10^{-3}$ PPO (baseline) $2.76\text{\times}{10}^{-01}$ $1.14\text{\times}{10}^{-01}$ $1.39\text{\times}{10}^{-01}$ $1.67\text{\times}{10}^{-01}$ $10^{-2}$ DPO (basic) $9.19\text{\times}{10}^{-03}$ $3.91\text{\times}{10}^{-03}$ $1.79\text{\times}{10}^{-03}$ $2.57\text{\times}{10}^{-03}$ $10^{-2}$ PG–DPO (basic) $5.57\text{\times}{10}^{-04}$ $1.09\text{\times}{10}^{-04}$ $9.20\text{\times}{10}^{-05}$ $6.00\text{\times}{10}^{-05}$ $10^{-2}$ DPO (distill.) $4.64\text{\times}{10}^{-03}$ $2.55\text{\times}{10}^{-03}$ $1.31\text{\times}{10}^{-03}$ $1.80\text{\times}{10}^{-03}$ $10^{-2}$ PG–DPO (distill.) $5.60\text{\times}{10}^{-04}$ $1.09\text{\times}{10}^{-04}$ $9.15\text{\times}{10}^{-05}$ $6.00\text{\times}{10}^{-05}$ $10^{-2}$ PPO (baseline) $3.00\text{\times}{10}^{-01}$ $7.88\text{\times}{10}^{-02}$ $1.50\text{\times}{10}^{-01}$ $1.58\text{\times}{10}^{-01}$ $10^{-1}$ DPO (basic) $1.04\text{\times}{10}^{-02}$ $4.21\text{\times}{10}^{-03}$ $1.81\text{\times}{10}^{-03}$ $2.50\text{\times}{10}^{-03}$ $10^{-1}$ PG–DPO (basic) $5.29\text{\times}{10}^{-03}$ $1.03\text{\times}{10}^{-03}$ $8.68\text{\times}{10}^{-04}$ $5.76\text{\times}{10}^{-04}$ $10^{-1}$ DPO (distill.) $6.23\text{\times}{10}^{-03}$ $2.76\text{\times}{10}^{-03}$ $1.33\text{\times}{10}^{-03}$ $1.53\text{\times}{10}^{-03}$ $10^{-1}$ PG–DPO (distill.) $5.33\text{\times}{10}^{-03}$ $1.03\text{\times}{10}^{-03}$ $8.73\text{\times}{10}^{-04}$ $5.76\text{\times}{10}^{-04}$ $10^{-1}$ PPO (baseline) $2.70\text{\times}{10}^{-01}$ $9.65\text{\times}{10}^{-02}$ $1.55\text{\times}{10}^{-01}$ $1.78\text{\times}{10}^{-01}$
Misaligned
$s$ Method $d=5$ $10$ $50$ $100$ $10^{-3}$ DPO (basic) $7.49\text{\times}{10}^{-03}$ $3.65\text{\times}{10}^{-03}$ $1.85\text{\times}{10}^{-03}$ $2.56\text{\times}{10}^{-03}$ $10^{-3}$ PG–DPO (basic) $1.94\text{\times}{10}^{-04}$ $1.55\text{\times}{10}^{-04}$ $4.30\text{\times}{10}^{-04}$ $8.89\text{\times}{10}^{-04}$ $10^{-3}$ DPO (distill.) $4.49\text{\times}{10}^{-03}$ $2.70\text{\times}{10}^{-03}$ $1.38\text{\times}{10}^{-03}$ $1.02\text{\times}{10}^{-03}$ $10^{-3}$ PG–DPO (distill.) $1.95\text{\times}{10}^{-04}$ $1.55\text{\times}{10}^{-04}$ $4.30\text{\times}{10}^{-04}$ $8.89\text{\times}{10}^{-04}$ $10^{-3}$ PPO (baseline) $2.87\text{\times}{10}^{-01}$ $8.25\text{\times}{10}^{-02}$ $1.51\text{\times}{10}^{-01}$ $1.69\text{\times}{10}^{-01}$ $10^{-2}$ DPO (basic) $7.87\text{\times}{10}^{-03}$ $3.50\text{\times}{10}^{-03}$ $2.11\text{\times}{10}^{-03}$ $2.68\text{\times}{10}^{-03}$ $10^{-2}$ PG–DPO (basic) $85\text{\times}{10}^{-03} 1.1$$19\text{\times}{10}^{-03}$ $2.07\text{\times}{10}^{-03}$ $2.62\text{\times}{10}^{-03}$ $10^{-2}$ DPO (distill.) $4.86\text{\times}{10}^{-03}$ $2.69\text{\times}{10}^{-03}$ $2.04\text{\times}{10}^{-03}$ $1.46\text{\times}{10}^{-03}$ $10^{-2}$ PG–DPO (distill.) $85\text{\times}{10}^{-03} 1.1$$19\text{\times}{10}^{-03}$ $2.07\text{\times}{10}^{-03}$ $2.62\text{\times}{10}^{-03}$ $10^{-2}$ PPO (baseline) $2.58\text{\times}{10}^{-01}$ $8.99\text{\times}{10}^{-02}$ $1.59\text{\times}{10}^{-01}$ $1.38\text{\times}{10}^{-01}$ $10^{-1}$ DPO (basic) $1.44\text{\times}{10}^{-01}$ $4.81\text{\times}{10}^{-01}$ $5.94\text{\times}{10}^{-01}$ $3.74\text{\times}{10}^{-01}$ $10^{-1}$ PG–DPO (basic) $2.40\text{\times}{10}^{-02}$ $1.33\text{\times}{10}^{-02}$ $2.29\text{\times}{10}^{-02}$ $9.31\text{\times}{10}^{-03}$ $10^{-1}$ DPO (distill.) $9.67\text{\times}{10}^{-03}$ $2.26\text{\times}{10}^{-01}$ $2.34\text{\times}{10}^{-01}$ $1.24\text{\times}{10}^{-01}$ $10^{-1}$ PG–DPO (distill.) $1.23\text{\times}{10}^{-02}$ $6.63\text{\times}{10}^{-03}$ $1.34\text{\times}{10}^{-02}$ $6.93\text{\times}{10}^{-03}$ $10^{-1}$ PPO (baseline) $2.78\text{\times}{10}^{-01}$ $8.84\text{\times}{10}^{-02}$ $1.47\text{\times}{10}^{-01}$ $1.77\text{\times}{10}^{-01}$
Table 1: Decision-time Euclidean RMSE at $t=0$ (tail median over the late-training window; last six evaluation snapshots), computed using the common Euclidean-norm RMSE definition (65). DPO rows report the deployable policy output (Stage 1). PG–DPO rows report the post-hoc Pontryagin projection output (Stage 2; residual form). “Distill.” rows correspond to the amortized deployable policy trained via interactive distillation. PPO is a model-free baseline trained under the same benchmark setting (same simulator, horizon $T=1.5$, and $\theta$-blind restriction). Bold entries indicate the best (lowest) RMSE among the five methods for each $(s,d)$ and geometry.
@@ 4.3 A strong RL baseline: PPO, and why it falls short in our benchmark
Since classical HJB solvers and value-function-based deep PDE surrogates are not practical baselines in the high-dimensional uncertain regime targeted here (Section C), PPO serves as a strong simulation-only comparator that operates on the same sampled trajectories without exploiting value-function PDE structure. For a fair comparison, PPO is trained on the same Euler simulator and time discretization as our DPO/PG–DPO setup, under the same deployability restriction (the policy never observes the latent $\theta$), and under the same terminal-utility objective. We match simulation budgets at the level of total environment interactions (simulated trajectories $\times$ time steps); implementation and PPO tuning details are deferred to the appendix and code release.
Empirical outcome. Table 1 shows that PPO remains far from the analytic decision-time $q$-reference across essentially all conditions, with RMSE typically on the order of $10^{-1}$. In contrast, the Pontryagin-based pipeline attains substantially smaller errors: in aligned regimes the Stage 2 projection (PG–DPO) reaches the $10^{-5}$$10^{-4}$ range for small and moderate uncertainty, while in misaligned regimes the projection advantage narrows but remains systematic. Distillation improves the deployable Stage 1 policy relative to basic Stage 1 (DPO), but does not eliminate the remaining gap to the post-hoc projection, consistent with the amortization interpretation in Section 4.2.
@@
In contrast, Stage 1 (DPO) exploits backpropagation through the differentiable simulator (pathwise gradients), and Stage 2 leverages the affine-in-control Pontryagin structure through a $q$-aggregated projection, replacing a noisy high-dimensional policy-gradient update by a structured stationarity correction tailored to the $\theta$-blind ex–ante objective.
@@ 5 Recovering Intertemporal Hedging Demand in Factor-Driven Markets
Section 4 stressed scaling under static drift uncertainty, where the target $q$-reference is time-homogeneous and largely myopic. Here we shift the focus to an economic target: recovering the intertemporal hedging demand induced by factor-driven investment opportunities when return shocks are correlated with factor shocks (Campbell and Viceira, 2002; Xia, 2001).
We adopt the mean-reverting Gaussian premium benchmark of Section 2.3.2. At decision time ($t=0$), the investor is assumed to have access only to an interval/posterior uncertainty description $q_{0}(dy)$ for the initial premium state $Y_{0}$ (e.g., $Y_{0}\sim\mathcal{N}(m_{0},P_{0})$), and we restrict attention to deployable $Y$-blind policies that depend only on observable wealth and time-to-go. When return and factor shocks are correlated ($\rho\neq 0$), the analytic OU decision-time reference (37) contains the cross term $M_{\mathrm{cross}}(\tau)$ (with $\tau:=T-t$) that encodes intertemporal hedging demand.
We evaluate: (i) Stage 1 DPO (a $Y$-blind warm-up trained by pathwise/BPTT gradients), (ii) Stage 2 Pontryagin projection completing PG–DPO (a post-hoc $q_{0}$-aggregated stationarity correction), and (iii) interactive distillation (amortization), against a model-free PPO baseline—all under the same $Y$-blind deployability restriction. Performance is measured by decision-time RMSE at $t=0$ relative to the analytic constant-portfolio OU reference (37).
@@ 5.1 Experimental setting
Decision-time interval/posterior estimate for $Y_{0}$ and deployability. Throughout this section, all reported policies are deployable and $Y$-blind: the control is a function of observable wealth and time-to-go only, and never takes as input (i) a point estimate of $Y_{0}$, (ii) the realized initial premium $Y_{0}$, or (iii) the factor path $(Y_{t})$ (including the PPO baseline). At deployment, the investor is assumed to have access only to an uncertainty description $q_{0}(dy)$ for $Y_{0}$ (interval/posterior output of an external estimation pipeline), and commits to a single policy optimized ex ante under $Y_{0}\sim q_{0}$ (without filtering/updating over $[0,T]$).
@@
Two-stage solver, amortization, and evaluation protocol. We use the two-stage pipeline of Section 3. Stage 1 (DPO) trains a deployable $Y$-blind policy by stochastic gradient ascent using pathwise/BPTT gradients (Section 3.1). Stage 2 (Pontryagin projection) applies the $q_{0}$-aggregated stationarity correction computed under a warm-up policy (Section 3.2), implemented in the residual/control-variate form (Section G.1); the resulting projected rule is the output of PG–DPO. Interactive distillation amortizes the projected teacher into a fast deployable policy network (Section G.2). As a model-free baseline, we also train a PPO policy under the same $Y$-blind observation restriction and report its decision-time full RMSE in Table 2.
@@
In addition to the full allocation error, we use the myopic+hedging decomposition induced by the OU structure, and report component-wise diagnostics for the projected (Stage 2 / PG–DPO) rules. To reduce noise from stochastic optimization, for each condition we summarize each metric by a tail median over the last six evaluation checkpoints.
@@ 5.2 Results: hedging-demand recovery, amortization, and robustness to decision-time uncertainty
Aligned $P_{0}$
$s_{0}$ Method $d=5$ $10$ $50$ $100$ $10^{-3}$ DPO (basic) $6.31\text{\times}{10}^{-03}$ $5.19\text{\times}{10}^{-03}$ $4.01\text{\times}{10}^{-03}$ $3.54\text{\times}{10}^{-03}$ $10^{-3}$ PG–DPO (basic) $4.71\text{\times}{10}^{-05}$ $5.10\text{\times}{10}^{-05}$ $1.39\text{\times}{10}^{-04}$ $1.56\text{\times}{10}^{-04}$ $10^{-3}$ DPO (distill.) $2.46\text{\times}{10}^{-03}$ $3.57\text{\times}{10}^{-03}$ $3.58\text{\times}{10}^{-03}$ $3.22\text{\times}{10}^{-03}$ $10^{-3}$ PG–DPO (distill.) $4.43\text{\times}{10}^{-05}$ $5.39\text{\times}{10}^{-05}$ $1.37\text{\times}{10}^{-04}$ $1.42\text{\times}{10}^{-04}$ $10^{-3}$ PPO (baseline) $7.78\text{\times}{10}^{-02}$ $1.03\text{\times}{10}^{-01}$ $4.20\text{\times}{10}^{00}$ $2.41\text{\times}{10}^{00}$ $10^{-2}$ DPO (basic) $5.61\text{\times}{10}^{-03}$ $4.83\text{\times}{10}^{-03}$ $3.83\text{\times}{10}^{-03}$ $3.50\text{\times}{10}^{-03}$ $10^{-2}$ PG–DPO (basic) $5.03\text{\times}{10}^{-05}$ $4.54\text{\times}{10}^{-05}$ $1.45\text{\times}{10}^{-04}$ $1.75\text{\times}{10}^{-04}$ $10^{-2}$ DPO (distill.) $3.08\text{\times}{10}^{-03}$ $3.16\text{\times}{10}^{-03}$ $3.53\text{\times}{10}^{-03}$ $3.21\text{\times}{10}^{-03}$ $10^{-2}$ PG–DPO (distill.) $4.48\text{\times}{10}^{-05}$ $5.11\text{\times}{10}^{-05}$ $1.47\text{\times}{10}^{-04}$ $1.56\text{\times}{10}^{-04}$ $10^{-2}$ PPO (baseline) $7.30\text{\times}{10}^{-02}$ $9.15\text{\times}{10}^{00}$ $4.27\text{\times}{10}^{00}$ $2.43\text{\times}{10}^{00}$ $10^{-1}$ DPO (basic) $6.93\text{\times}{10}^{-03}$ $4.53\text{\times}{10}^{-03}$ $3.96\text{\times}{10}^{-03}$ $3.47\text{\times}{10}^{-03}$ $10^{-1}$ PG–DPO (basic) $5.50\text{\times}{10}^{-05}$ $4.16\text{\times}{10}^{-05}$ $2.63\text{\times}{10}^{-04}$ $2.97\text{\times}{10}^{-04}$ $10^{-1}$ DPO (distill.) $2.85\text{\times}{10}^{-03}$ $3.70\text{\times}{10}^{-03}$ $3.65\text{\times}{10}^{-03}$ $3.28\text{\times}{10}^{-03}$ $10^{-1}$ PG–DPO (distill.) $4.77\text{\times}{10}^{-05}$ $5.50\text{\times}{10}^{-05}$ $2.60\text{\times}{10}^{-04}$ $2.88\text{\times}{10}^{-04}$ $10^{-1}$ PPO (baseline) $7.34\text{\times}{10}^{-02}$ $1.00\text{\times}{10}^{-01}$ $4.22\text{\times}{10}^{00}$ $2.64\text{\times}{10}^{00}$
Misaligned $P_{0}$
$s_{0}$ Method $d=5$ $10$ $50$ $100$ $10^{-3}$ DPO (basic) $6.11\text{\times}{10}^{-03}$ $5.40\text{\times}{10}^{-03}$ $3.87\text{\times}{10}^{-03}$ $3.62\text{\times}{10}^{-03}$ $10^{-3}$ PG–DPO (basic) $5.12\text{\times}{10}^{-05}$ $5.11\text{\times}{10}^{-05}$ $1.35\text{\times}{10}^{-04}$ $1.57\text{\times}{10}^{-04}$ $10^{-3}$ DPO (distill.) $3.51\text{\times}{10}^{-03}$ $2.93\text{\times}{10}^{-03}$ $3.56\text{\times}{10}^{-03}$ $3.21\text{\times}{10}^{-03}$ $10^{-3}$ PG–DPO (distill.) $4.36\text{\times}{10}^{-05}$ $5.48\text{\times}{10}^{-05}$ $1\text{\times}{10}^{-04} 1.44×1$$4\text{\times}{10}^{-04}$ $10^{-3}$ PPO (baseline) $8.96\text{\times}{10}^{-02}$ $9.28\text{\times}{10}^{-02}$ $4.37\text{\times}{10}^{00}$ $2.46\text{\times}{10}^{00}$ $10^{-2}$ DPO (basic) $6.17\text{\times}{10}^{-03}$ $4.89\text{\times}{10}^{-03}$ $3.89\text{\times}{10}^{-03}$ $3.57\text{\times}{10}^{-03}$ $10^{-2}$ PG–DPO (basic) $4.99\text{\times}{10}^{-05}$ $5.96\text{\times}{10}^{-05}$ $1.54\text{\times}{10}^{-04}$ $1.75\text{\times}{10}^{-04}$ $10^{-2}$ DPO (distill.) $2.84\text{\times}{10}^{-03}$ $3.97\text{\times}{10}^{-03}$ $3.68\text{\times}{10}^{-03}$ $3.21\text{\times}{10}^{-03}$ $10^{-2}$ PG–DPO (distill.) $4.63\text{\times}{10}^{-05}$ $6.31\text{\times}{10}^{-05}$ $1.54\text{\times}{10}^{-04}$ $1.57\text{\times}{10}^{-04}$ $10^{-2}$ PPO (baseline) $8.02\text{\times}{10}^{-02}$ $7.75\text{\times}{10}^{00}$ $4.47\text{\times}{10}^{00}$ $2.53\text{\times}{10}^{00}$ $10^{-1}$ DPO (basic) $6.09\text{\times}{10}^{-03}$ $4.17\text{\times}{10}^{-03}$ $3.97\text{\times}{10}^{-03}$ $3.56\text{\times}{10}^{-03}$ $10^{-1}$ PG–DPO (basic) $5.63\text{\times}{10}^{-05}$ $2.44\text{\times}{10}^{-04}$ $3.24\text{\times}{10}^{-04}$ $3.29\text{\times}{10}^{-04}$ $10^{-1}$ DPO (distill.) $3.22\text{\times}{10}^{-03}$ $3.08\text{\times}{10}^{-03}$ $3.65\text{\times}{10}^{-03}$ $3.18\text{\times}{10}^{-03}$ $10^{-1}$ PG–DPO (distill.) $5.31\text{\times}{10}^{-05}$ $2.46\text{\times}{10}^{-04}$ $3.22\text{\times}{10}^{-04}$ $3.21\text{\times}{10}^{-04}$ $10^{-1}$ PPO (baseline) $6.30\text{\times}{10}^{-02}$ $1.08\text{\times}{10}^{-01}$ $4.26\text{\times}{10}^{00}$ $2.46\text{\times}{10}^{00}$
Table 2: Decision-time RMSE at $t=0$ in the OU premium benchmark (tail median over the last six evaluation checkpoints), sweeping the decision-time uncertainty scale $s_{0}$ in $Y_{0}\sim\mathcal{N}(m_{0},P_{0})$ under both aligned (66) and misaligned (67) geometries. DPO rows report the Stage 1 warm-up (deployable) policy output. PG–DPO rows report the Stage 2 post-hoc $q_{0}$-aggregated Pontryagin projection output (residual/control-variate implementation; Section G.1) computed under the corresponding warm policy. “Distill.” rows use interactive distillation (Section G.2) to amortize the projected teacher into a deployable network. RMSE is computed identically for all methods (no method-specific or post-hoc scale correction). The PPO baseline is frequently unstable in this terminal-only, $Y$-blind setting; very large entries (e.g., $s_{0}=10^{-2},\,d=10$) reflect genuine training divergence rather than typographical error.
Metrics and diagnostics. Table 2 reports the full decision-time RMSE at $t=0$ for all deployable policies, including the PPO baseline. To isolate the economic hedging channel, we use the OU-induced decomposition of the analytic reference
@@
where $\pi^{\mathrm{myo}}$ is the independence-case reference obtained by setting $\rho=0$ (equivalently, dropping $M_{\mathrm{cross}}(\tau)$ in (37)), and $\pi^{\mathrm{hedge}}:=\pi^{\mathrm{ref}}-\pi^{\mathrm{myo}}$ is the residual $\rho$-dependent term. We report the hedging-component RMSE for the projected (Stage 2 / PG–DPO) rules in Table 3; the myopic-component RMSE and hedging-direction cosine similarity are deferred to Appendix I. Since Stage 1 (DPO) and PPO do not expose a compatible myopic/hedging split under our diagnostic protocol, component-wise diagnostics are reported only for the projected rules.